OpenStudy (anonymous):

A motorcycle is moving in a rectilinear road AB, lefting from A with the constant velocity v1=20km/h. When it reaches the midpoint M of AB, the velocity abruptly changes to v2=80km/h and it keeps constant until it reaches point B. Find the average scalar velocity of all the way. Pls help! Answer is 32km/h but I don't get why.

OpenStudy (anonymous):

Theres none

OpenStudy (mani_jha):

Average velocity = Total Displacement/ Total Time taken Let the length of AB =x km Time taken to cover first half =(x/2)/20=x/40 Time taken to cover second half = (x/2)//80=x/160 Total displacement =x Total time = x/40+x/160=x/32 Thus, average velocity =x/(x/32) =32 km/hr

OpenStudy (ash2326):

Let's assume the distance to be x km For x/2 km its speed is 20 km/h time taken=\[\frac{\frac{x}{2}}{20}=\frac{x}{40} hours\] For other half, its speed is 80kmph so time taken=\[\frac{\frac{x}{2}}{80}=\frac{x}{160} hours\] Total time taken= \[\frac{x}{40}+\frac{x}{160}=\frac{5x}{160}\] Average Speed = distance/total time taken=\[\frac{x}{\frac{5x}{160}}=\frac{160}{5x}=32km/h\]

OpenStudy (anonymous):

Thanks guys, you're amazing!!